OpenStudy (anonymous):

f(x) = (x-3)/8 ; g(x) = 8x + 3, find g(f(x)). I got g(f(x)) = x - (3/8) is this correct?

OpenStudy (anonymous):

g(x) = 8x + 3 g((x-3)/8 ) = 8((x-3)/8 ) + 3 g((x-3)/8 ) = x